1.(p∧q)=(q∨p) and (p∨q)=(q∨p)
implies an _____
a.Associative laws
b.Distributive Laws
c.Commutative Laws
d.Idempotent Laws

2.____is equivalent to (p∨q)
a.(p↔q)
b.(∼p∧q)
c.(∼∼q)
d.∼(∼p∧∼q)

3.The_____ of a matrix is the inter- changing of its row with the column
a.diagonal matrix
b.zero matrix
c.transpose matrix
d.identity matrix

4.A matrix that has elements only on its diagonal is called
a.transpose matrix
b.diagonal matrix
c.identity matrix
d.zero matrix

5.given that A= ⎛1 2 3⎞
⎝4 5 6⎠ and


⎛1 2⎞
B= ⎜3 4⎟
⎝5 6⎠



. Find AB

a.⎛ 0 12 17⎞
⎜ 7 10 31⎟
⎝20 40 45⎠


b.⎛ 0 12 17⎞
⎜ 19 26 31⎟
⎝ 20 40 45⎠

c.⎛ 0 12 17⎞
⎜ 19 26 31⎟
⎝ 29 40 51⎠

d.⎛ 5 12 15⎞
⎜ 19 26 31⎟
⎝ 29 40 51⎠
1
Expert's answer
2019-03-13T11:17:28-0400

1) (p∧q)=(q∧p) and (p∨q)=(q∨p) implies a commutative law (c)


2)∼(∼p∧∼q) is equivalent to (p∨q)  (d)


3)The transpose matrix of a matrix is the inter- changing of its row with the column (a)


4)A matrix that has elements only on its diagonal is called diagonal matrix (b)


5)

"\begin{pmatrix}
 1 & 2 & 3 \\ 
 4 & 5 & 6

\end{pmatrix}
\begin{pmatrix}
 1 & 2 \\
 3 & 4 \\
 5 & 6
\end{pmatrix}
=
\begin{pmatrix}
 22 & 29 \\
 49 & 64 \\
 
\end{pmatrix}"


No correct answer is given